An exam consists of 50 multiple choice questions. Based on how much you studied, for any given question, you think you have a probability of 0.64 of getting the correct answer. Consider the sampling distribution of the sample proportion of correct questions out of 50.
Find the mean and standard error of the sampling distribution of this proportion.

Answer by stanbon(57219)

An exam consists of 50 multiple choice questions. Based on how much you studied, for any given question, you think you have a probability of 0.64 of getting the correct answer. Consider the sampling distribution of the sample proportion of correct questions out of 50.
Find the mean and standard error of the sampling distribution of this proportion.
-------------
Binomial problem with n = 50 and p = 0.64
mean = np = 50*0.64
std = sqrt(npq) = sqrt(50*0.64*0.36)
==========================================
Cheers,
Stan H.